ganeshie8 (ganeshie8):

show that \((n-2)!\) is divisible by \(n-1\) when ever \(n \ne p+1\) where \(p\) is a prime and \(n\gt 5\)

OpenStudy (bubblegum.):

i took n=5 so \(\Large\frac{(5-2)!}{5-1}\) = \(\Large\frac{3 \times 2 \times 1}{4}\)=\(\Large\frac{3}{2}\) so here (n-2)! is not divisible by (n-1)

OpenStudy (anonymous):

hm

ganeshie8 (ganeshie8):

corrected the question, sorry

ganeshie8 (ganeshie8):

@bubblegum.

OpenStudy (bubblegum.):

\(\large n \ne p+1\) \(\large (n-1) \ne p\) we can write \(\large\frac{(n-2)!}{(n-1)}\) as \(\large\frac{(n-1)!}{(n-1)^2}\) If we have any number say X lets take X=36 then the divisors are->1,2,3,4,6,9,12,18,36 we can write 36 as a product of 2 divisors like-> 1x36=36 2x18=36 3x12=36... similarly for (n-1) the divisors will obviously be smaller than (n-1) and they all will come in the factorial we know that every number X can be represented as a product of its 2 factors lets say \((n-1)=d_1 \times d_2\) where \(d_1\) and \(d_2\) are the 2 divisors of \(n-1\) multiplying them in numerator gives us (n-1) we will have \((n-1)!=(n-1)(n-2)(n-3)...d_1...d_2..2.1\) we can see we have \((n-1)^2\) term generated in the numerator and we can easily say that this numerator will be divisible by denominator which is \((n-1)^2\)

OpenStudy (bubblegum.):

we take \(n > 5\) because we cannot represent 1,2,3,4,5 is the form of product of 2 divisors now m thing why \(n \ne p+1\)

OpenStudy (bubblegum.):

*thinking

OpenStudy (bubblegum.):

if \(n=1+p\) then \(n-1=p\) \((n-1)!=p!\) yeah its clear that p! will obviously not contain \(p^2\) term so clearly \((n-1)!\) will not be divisible by \((n-1)^2\)

ganeshie8 (ganeshie8):

Very nice!

OpenStudy (bubblegum.):

thanks :)